# from perception to punchline: empowering vlm with the art of in-the-wild meme

### v1

@@ 5.1 Meme Quality and Diversity with HUMOR
We evaluate the proposed HUMOR framework against several competitive baselines and model variants. Concretely, the compared systems include multiple open-source and closed-source VLMs, as well as our HUMOR-CoT model, which is fine-tuned with the hierarchical CoT design. To further investigate the efficacy of alternative CoT methods for meme generation, we also include several advanced CoT frameworks (kim2023cot; chen2024huatuogpt), all trained under the same data and protocol. See Appendix  H for detailed training settings and Appendix G for the details of datasets. Given the highly open-ended and human-aligned nature of meme generation, we prioritize human evaluation. Human annotators are asked to assign scores to generated memes along four predefined quality axes. In addition, we adopt the conventional metric of text-level similarity between generated captions and their original reference texts. To further quantify generation diversity, we introduce a novel metric called Distance under Context Swap. This measure replaces the original context in the training set with a randomly selected one (kept consistent across models), and computes the textual divergence between the resulting caption and the original. A larger distance suggests reduced overfitting to SFT labels and better adaptability to new contexts. Due to observed instability in VLM-based rubric scores for meme evaluation (Sec. 5.2), we incorporate only one VLM-based metric: a human-likeness score. This is formulated as a binary classification estimating the probability that a meme was created by a human, with higher values indicating better. We adopt Gemini-2.5-pro as the evaluator for computing Human Rate, as it demonstrates the most stable and consistent behavior among candidate VLMs in our evaluator reliability analysis (Appendix  J). For a more detailed description of the indicator meanings and evaluation criteria, see Appendix  I.
@@
Qwen2.5-7B-Instruct (bai2025qwen2) 2.39 3.35 2.91 2.57 0.576 0.564 75.7
Qwen2.5-32B-Instruct (bai2025qwen2) 2.54 3.52 3.09 2.76 0.564 0.566 82.2
InternVL3-8B (zhu2025internvl3) 2.39 2.79 3.04 2.79 0.545 0.564 62.7
GLM-4.1V-9B-Thinking (hong2025glm) 1.73 2.62 2.75 2.71 0.602 0.572 45.1
Keye-VL-8B-preview (team2025kwai) 2.35 3.19 2.99 2.71 0.585 0.580 69.0
@@
GPT-4o (openai_gpt4o_blog_2024) 2.70 2.99 3.21 2.97 0.603 0.552 91.3
Gemini-2.5-flash (comanici2025gemini) 2.81 3.29 3.25 2.88 0.600 0.561 -
@@
   CoT with Single Path (kim2023cot) 1.87 2.79 2.68 2.45 0.637 0.570 86.0
   CoT with Self-Improve (chen2024huatuogpt) 2.38 3.68 3.00 2.65 0.629 0.578 89.1
   CoT with Subquestion (wei2022chain) 1.85 3.32 2.58 2.47 0.639 0.597 87.2
@@ Appendix A Hierarchical Chain-of-Thoughts of Metaphor
We collected over 4,000 meme images from public dataset (xu2024generating), and established a multi-dimensional labeling system:
@@ I.2 MaxDiff Ordering
Maximum Difference Scaling (MaxDiff), also known as best–worst scaling, is a widely used method in marketing science and preference elicitation louviere1991best; louviere2015best. In a typical MaxDiff task, respondents are repeatedly presented with small subsets of items (e.g., 3–5 candidates) and asked to indicate which option they consider the “best” and which the “worst.” Compared to traditional rating scales, MaxDiff provides more discriminative and reliable preference estimates because each choice yields two pieces of information: a positive preference for the selected “best” item and a negative preference for the “worst.”
The required number of tasks in MaxDiff depends on the total number of items $J$ to be evaluated and the subset size $k$. A common guideline is that each item should appear across multiple choice sets to ensure stable estimation. For example, using balanced incomplete block designs (BIBD), each respondent typically completes between $\frac{3J}{k}$ and $\frac{5J}{k}$ choice tasks to achieve acceptable reliability orme2010getting. Thus, the total number of questions can be determined systematically to balance respondent burden and statistical efficiency.

### v2

@@ 5.1 Meme Quality and Diversity with HUMOR
We evaluate the proposed HUMOR framework against several competitive baselines and model variants. Concretely, the compared systems include multiple open-source and closed-source VLMs, as well as our HUMOR-CoT model, which is fine-tuned with the hierarchical CoT design. To further investigate the efficacy of alternative CoT methods for meme generation, we also include several advanced CoT frameworks (Kim et al., 2023; Chen et al., 2024a), all trained under the same data and protocol. See Appendix  H for detailed training settings and Appendix G for the details of datasets. Given the highly open-ended and human-aligned nature of meme generation, we prioritize human evaluation. Human annotators are asked to assign scores to generated memes along four predefined quality axes. In addition, we adopt the conventional metric of text-level similarity between generated captions and their original reference texts. To further quantify generation diversity, we introduce a novel metric called Distance under Context Swap. This measure replaces the original context in the training set with a randomly selected one (kept consistent across models), and computes the textual divergence between the resulting caption and the original. A larger distance suggests reduced overfitting to SFT labels and better adaptability to new contexts. Due to observed instability in VLM-based rubric scores for meme evaluation (Sec. 5.2), we incorporate only one VLM-based metric: a human-likeness score. This is formulated as a binary classification estimating the probability that a meme was created by a human, with higher values indicating better. We adopt Gemini-2.5-pro as the evaluator for computing Human Rate, as it demonstrates the most stable and consistent behavior among candidate VLMs in our evaluator reliability analysis (Appendix  J). For a more detailed description of the indicator meanings and evaluation criteria, see Appendix  I.
@@
Qwen2.5-7B-Instruct (Bai et al., 2025) 2.39 3.35 2.91 2.57 0.576 0.564 75.7
Qwen2.5-32B-Instruct (Bai et al., 2025) 2.54 3.52 3.09 2.76 0.564 0.566 82.2
InternVL3-8B (Zhu et al., 2025b) 2.39 2.79 3.04 2.79 0.545 0.564 62.7
GLM-4.1V-9B-Thinking (Hong et al., 2025) 1.73 2.62 2.75 2.71 0.602 0.572 45.1
Keye-VL-8B-preview (Team et al., 2025) 2.35 3.19 2.99 2.71 0.585 0.580 69.0
@@
GPT-4o (OpenAI, 2024) 2.70 2.99 3.21 2.97 0.603 0.552 91.3
Gemini-2.5-flash (Comanici et al., 2025) 2.81 3.29 3.25 2.88 0.600 0.561 -
@@
   CoT with Single Path (Kim et al., 2023) 1.87 2.79 2.68 2.45 0.637 0.570 86.0
   CoT with Self-Improve (Chen et al., 2024a) 2.38 3.68 3.00 2.65 0.629 0.578 89.1
   CoT with Subquestion (Wei et al., 2022) 1.85 3.32 2.58 2.47 0.639 0.597 87.2
@@ Appendix A Hierarchical Chain-of-Thoughts of Metaphor
We collected over 4,000 meme images from public dataset (Xu et al., 2024), and established a multi-dimensional labeling system:
@@ I.2 MaxDiff Ordering
Maximum Difference Scaling (MaxDiff), also known as best–worst scaling, is a widely used method in marketing science and preference elicitation Louviere and Woodworth (1991); Louviere et al. (2015). In a typical MaxDiff task, respondents are repeatedly presented with small subsets of items (e.g., 3–5 candidates) and asked to indicate which option they consider the “best” and which the “worst.” Compared to traditional rating scales, MaxDiff provides more discriminative and reliable preference estimates because each choice yields two pieces of information: a positive preference for the selected “best” item and a negative preference for the “worst.”
The required number of tasks in MaxDiff depends on the total number of items $J$ to be evaluated and the subset size $k$. A common guideline is that each item should appear across multiple choice sets to ensure stable estimation. For example, using balanced incomplete block designs (BIBD), each respondent typically completes between $\frac{3J}{k}$ and $\frac{5J}{k}$ choice tasks to achieve acceptable reliability Orme (2010). Thus, the total number of questions can be determined systematically to balance respondent burden and statistical efficiency.